CAT Chat Minutes

February 28, 2005

 

Announcements and Follow-up
Mail Stops – Gwendolyn Morrison, PFS, has agreed to look at the problem of insufficient mailstops and will report back.

Long Range Schedule – The proposed high-current period, from August 9 at 8 a.m. to August 15, 2005 at 4 p.m., has been canceled.

Users wanted to know if the high-current period in August will now become a 100 mA low- emittance top-up period.

The decision has not been made at this time.   The long range operations schedule will be revised as appropriate.

Users are not satisfied with the current long range schedule wherein beam is shutdown for eight-hour periods from midnight to 8 am.   Some users would like to see a 24-hour shutdown.   They wonder why the eight-hour shutdowns are even necessary.

It was explained that it was user dissatisfaction with the previous 8 a.m. to 4 p.m. shutdown periods that prompted the current eight-hour period scheduling.   It was agreed that APS management will look into this issue.

Machine Stats
Delivered beam:  587.08 hrs.
Percentage of scheduled time:   97.85%
Downtime during period:   12.88 hrs.
Mean time between faults:   83.87 hrs.

Machine Faults, 2/22-28/2005:

Fill

Reason

Length of Downtime

#11   02/22 08:00 to 02/28 01:18   137.30 hours

Power bump ( ComEd)

6.70 hours, power event, affected inj. & SR

Beamline Data
Beamtime downtime summary by APS groups, 02/18-24/2005

DIN

Date

Beamline

Status

Duration

Group

Description

Resolution

501

02/19/05

2-ID

Closed

33 min

AOD-IT

S. Vogt at 2-ID-E reported loss of access to PV's, the " caget" and "caput" commands are not working.

S. Vogt reported that the PV's were working. ASD-IT and M. Smith looking into why this occurred; Smith believes that the gateway may have rebooted itself.

504

02/23/05

33-BM

Closed

3 hr 27 min

ASD-SI

33-BM-B shutter will not open. 33-BM-B door will not open.

It was determined that a simultaneous lock command (from a completed Search & Secure) and an unlock command (from depressing the pneumatic door open button) inhibited the function of the shutter and the door. Events caused timing problem within the PSS logic which failed "safe" by not allowing operation of the shutter, nor access to the hutch. Timing problem corrected, tested, and beamline restored to operations.

505

02/23/05

11-ID

Closed

2 hr 40 min

ASD-SI

Fault #395 A_FLOW_TO_0, Fault #67 PSS_NO_SWITCH-S.

The monitored flow on A_WBSA_F, B_WBSA_F were okay but the Love controllers would not reset.

ASD-SI replaced faulty Love Controllers. 1-for-1 replacement procedures followed, hardware tested. Beamline restored to proper operations.

Safety
The DOE Office of Independent Assessment is coming again in April, 2005, to observe: CAT safety plan, integrated user safety orientation and follow-thru with user training, CNM construction safety, and beryllium safety.

Remember to follow the guidelines posted on the web for safe battery recycling and tape used battery terminals.  

March 2005 safety inspections are scheduled for buildings 434 and 435.